OpenStudy (anonymous):

Write the equation of the line that passes through (2, 3) and is parallel to the line 12x – 5y = 2.

OpenStudy (anonymous):

Solve the equation for y.\[y=\frac{12 x}{5}-\frac{2}{5} \] The slope is the coefficent of x.\[\frac{y-3}{x-2}=\frac{12}{5} \]Solve the above for y.\[y=\frac{12 x}{5}-\frac{9}{5} \]

OpenStudy (anonymous):

thanks !

OpenStudy (anonymous):

Thank you for the medal.